Let X and Y ~U(0, 1]. X and Y are independent

a) Find the PDF of X+Y

b) Suppose now X~(0, a] Y~(0,b] and Ο <α<b. Find the PDF of X+Y
Answer #1

i have done if . firstly i answered the question 1 and then question 2, please go through the images for the solutions of the problems .

ut, x and y muco,t) .X, Y independent, a. 2 = x+y, let, Now, fx(x) = fy (Y) = { 1 if ocard loo.w. rify(@)= a, acxcb.) when, a

Now, max name of 2Exty will be a [when, x=1, 4=t]. and minimum value will be 22.05 when x=0, y =o) Now, for OcZ 22, we will cVCOV & (2) = { z if olaLL a-z if LOLZ L2 to other wise 2) if anuco, as you v co b] - then, fx(x) = f 40 La La o o.W. and, fyfor, olz-yla. z-aLYLZ. Now, for, olzla f₂ (2) = f f (2-4) dy for alzab, f (2) = 1 fez-y) dy max Ca, b-a) for, blzzatb, 22 (2). The pdf of zis. f(2)=( f (z-g) ay if ocala f f(2-y) ay if azzub maxca b-a) .. 2b it blzL atb O O.W.